From GlueXWiki
Jump to: navigation, search

March 10, 2016 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. NIM Papers
  2. Detector Status
  3. Electronics (Fernando, Chris, Nick)
  4. fADC125 meeting (Naomi, Cody, Beni)
  5. Other


Participants: Naomi, Mike (CMU) Eugene, Luke, Beni, Chris, Simon, Nick, and Lubomir (JLab).

NIM papers

- Beni started working on the Introduction and Conclusion for the CDC paper.

- Lubomir is finalizing the 4-page NIM paper (attached) - a contribution to the VCI conference.

Detector Status

- Beni looked at the connector that was showing problems before (I22 at the lower outer side of the CDC), but couldn't find anything suspicious. Right now there are no problems with these channels. Beni set the CDC HV to 500V during the accelerator down time.

- Lubomir discussed the measurements with empty target (chart attached). When the target was emptied the currents in the central FDC channels went down from 0.6 to 0.2 uA. After 20 min. we removed the profiler (and the TPOL converter) and when the beam came back the currents went up to 0.4 uA. Plotted is also the current in the corrector magnet and it also shows some change (possibly due to the effect of the profiler?) but it is not convincing that this caused the change in the FDC current. Eugene suggested another explanation: the profiler is stopping some low energy photons that are otherwise producing high currents in the FDC. These hypotheses have to be checked with empty target again by inserting the profiler and turning off the slow feedback.

fADC125 firmware

From Naomi:

Cody is working on eliminating the timing errors from the firmware. Naomi has been testing the new versions using the prototype at CMU. A few differences in methodology have been straightened out but the timing errors are still being worked on. Beni said that it should be possible to use the Hall D crates next week.